OpenStudy (anonymous):

-4x - 5 > -9 and describe the graph of the solution. someone help

OpenStudy (anonymous):

First... this is an absolute value function/problem so you need to realize that there are essentially two sides to this equation. LEFT SIDE (NEGATIVE) 5x - 2 = -8 solve normally so add the 2 -8 + 2 = -6 so now you have 5x = -6 divide -6 by 5 to get x value x = -6/5 RIGHT SIDE (POSITIVE) 5x - 2 = 8 Add the 2 5x = 10 Divide by 5 to get x.... 10/5 = x In simplified terms x = 2 SO X CAN BE EITHER 2 or -6/5 I'm sure you can understand the rest?

OpenStudy (anonymous):

so is it? x > 1 open circle on 1, shading to the right